CARME currently in vacuum, pressure ~3E-10 mbar. Chamber to be vented tomorrow afternoon by vacuum division. We have active both pirani and hot cathode gauge to aid.
Equipment removed from the chamber in preparation for dismount of CARME from the ring this week. Equipment removed includes:
- Neg pump cabling
- Heating wire
- FEE electronic units and associated cabling (power, hdmi etc)
- Stainless steel FEE holders
- Red and blue cooling hoses
Attached image shows chamber following equipment removal. Adaptor cards still attached to feed-through flanges. ERNI connnectors examined, all fine except for FEE aida03 which has a bent pin (see attached image). Will dismount adaptor cards to examine further tomorrow. Feed-through flanges will also be examined to check for bent pins following adaptor card removal.
New CARME frame is part constructed and is in the CRYRING area. Requires rails and castors which should arrive this week. The height of the frame will require shortening, to be level with the current frame. This will be looked into further tomorrow. |